# A priori error analysis of the proximal Galerkin method

### v1

@@ 4. Application I: The obstacle problem
and the obstacle $\phi\in H^{1}(\Omega)\cap L^{\infty}(\Omega)$ is fixed. For simplicity, we assume that $\phi|_{\partial\Omega}=-\delta<0$ for some positive constant $\delta$. It is well-known that there exists a unique solution to 95 [18, Theorem 6.9-1], here denoted by $u^{\ast}$.
@@
[…]
@@
Main goal: We derive error estimates in Corollary 2 for the PG method applied to the obstacle problem (Algorithm 2) for the two choices of $V_{h}\times W_{h}$ given in 102 and 103, below. To this end, we utilize Theorem 3.8 and verify that Assumptions 2 and 3 hold for each pair of spaces, respectively. We follow with preliminary results for each case.
[…]
@@
(102) $V_{h}$ =$=\{v\in L^{\infty}(\Omega)\mid v|_{T}\in\mathbb{P}_{1}(T)\oplus\mathbb{B}(T)~{}\text{ for all }T\in\mathcal{T}_{h}\}\cap H^{1}_{0}(\Omega),$
@@
Both of the $V_{h}$ subspaces in 102 and 103 imply fixing the degrees of freedom of $u_{h}^{k}$ on the Dirichlet boundary. This is a textbook procedure that is simplified in our setting because of the homogeneous boundary conditions specified in 95. Likewise, the continuous Lagrange elements defining $W_{h}$ in 103 also require fixing the degrees of freedom of the latent variable $\psi_{h}^{k}$ on the Dirichlet boundary. In this case, we construct the lift
@@
where $\mathcal{N}_{h}\cap\partial\Omega$ are the boundary dofs and $\{\varphi_{z}\}_{z\in\mathcal{N}_{h}}$ are the global shape functions corresponding to the (nodal) dofs $\mathcal{N}_{h}$. This construction ensures that the bound-preserving discrete solution $\tilde{u}_{h}^{k}$ 100 satisfies the same homogeneous Dirichlet boundary conditions as the solution $u^{\ast}$ it’s meant to approximate. Indeed, observe that
@@
with $\nabla\mathcal{R}^{*}(\psi)=\exp(\psi)+\phi$ and $\nabla\mathcal{R}(u)=\ln(u-\phi)$ for the particular choice of Legendre function in Algorithm 2. In turn, when using the elements in Case II, we understand 99b as
@@
[…]
We begin with the Fortin operator for the subspaces in 102, and establish the operator’s approximation properties required in Assumption 3.
@@
The subspaces given in 102 satisfies the inf-sup condition 40 and there exists a stable Fortin operator satisfying 42 and
@@
where ~$\tilde{\mathcal{I}}_{h}:L^{1}(\Omega)\rightarrow\mathbb{P}_{1}(\mathcal{T}_{h})\cap H^{1}_{0}(\Omega)$ is the quasi-interpolant introduced in [24, Section 6], $I$ is the identity operator, and $\tilde{\Pi}_{h}:L^{2}(\Omega)\rightarrow V_{h}$ is defined element-wise to satisfy $(\tilde{\Pi}_{h}v)|_{T}:=b_{T}v_{T}\in\mathbb{B}(T)$, where $v_{T}\in\mathbb{P}_{0}(T)=\mathbb{R}$ solves
@@
Using 110 in 107 along with the triangle inequality and the stability of ~$\tilde{\mathcal{I}}_{h}$, we obtain the stability of $\Pi_{h}$ in $L^{2}(\Omega)$. To show stability in the $H^{1}$-semi norm, we apply triangle inequality, a local inverse estimate, 110 and the stability and approximation properties of ~$\tilde{\mathcal{I}}_{h}$ [24, Lemma 6.3 and Theorem 6.4]:
@@
where $\Delta_{T}$ is a macro-element. Summing over mesh elements and using mesh regularity shows stability in the $H^{1}(\Omega)$-seminorm. We conclude that 105 holds. This stability estimate and the observation that
@@
Note that the discrete iterates $u_{h}^{k}$ have bound-preserving local averages. Indeed, testing 99b by the indicator function of one element (an admissible test function because $W_{h}=\mathbb{P}_{0}(\mathcal{T}_{h})$) and using that $\nabla\mathcal{R}^{*}(\psi)>\phi$ shows that $u_{h}^{k}\in K_{h}$ for all $k$, with
@@
To satisfy Assumption 2, we first require a map defined over $V$ and mapping the set $K_{h}$ to the set $K$. We will later shift this map to construct the enriching map $\mathcal{E}_{h}$, see 129. The natural choice is the Clément interpolant [19, 5]. Here, we utilize a specialized variant of this interpolant that employs weighted local averages over vertex patches to maintain second-order accuracy for smooth functions. This specific interpolant was first introduced in [29], but we slightly modify its definition to incorporate non-homogeneous boundary data.
@@
[…]
@@
In the above, we also used that $\mathrm{card}(\omega_{z})$ is uniformly bounded w.r.t. $h$ for all $z\in\mathcal{N}_{h}$, which follows from the shape-regularity of the mesh; see, e.g., [25, Proposition 11.6]. If $z\in\partial\Omega$, then $z\in\partial T\cap\omega_{z}$ for some $T\subset\omega_{z}$. Using the definition of $\mathcal{C}_{h}$ on boundary nodes, we apply local inverse and trace inequalities to bound
@@
Combining 117 with 120 and using the Bramble–Hilbert Lemma (see, e.g., [24, Lemma 5.6]) yields the required bound on the first term of 115. To obtain the required bound on the second term, we apply triangle and inverse estimates:
@@ 4.2. Preliminaries for the $(\mathbb{P}_{1},\mathbb{P}_{1})$-element pair
For this case, the Fortin operator is simply the $L^{2}(\Omega)$-projection; cf. [36, Remark 5.2]. Thus, the main task is to construct the enriching map of Assumption 2, which will again depend on a modified Clément quasi-interpolant. Using the definition of $K_{h}$ 62 and testing with $\varphi_{z}$, we observe that
@@
If $z\in\partial\Omega$, we proceed as before and select $v_{z}=(\mathcal{SZ}_{h}v)(z)$. For simplicity, we make the following symmetry assumption on the mesh, which is sufficient for optimality of the quasi-interpolant given in 122. Note that this assumption is the same condition necessary for optimality of the classical Clément quasi-interpolant [29].
@@
We now prove that this condition implies optimality of 122.
@@
[…]
@@
The proof follows similar arguments to the proof of Lemma 4.2. We only highlight the key points. We first show that under 123, $\mathcal{C}_{h}p(z)=p(z)$ for all $p\in\mathbb{P}_{1}(\omega_{z})$. To see this first note that for linear $p$, we have
@@
where we used that $\int_{T}\varphi_{z^{\prime}}\varphi_{z}\operatorname{d}\!x=(n+1)^{-1}(n+2)^{-1}|T|(1+\delta_{z^{\prime},z})$ [17, Exercise 4.1.1]. Hence,
@@
In the last step, we used that $\int_{\omega_{z}}\varphi_{z}=|\omega_{z}|/(n+1)$ and 123. Along with the observation that $\overline{b}_{z}=b$ for any constant $b$, we conclude that $\mathcal{C}_{h}p=p$ for any $p\in\mathbb{P}_{1}(\omega_{z})$. In turn, it suffices to bound $\|\mathcal{C}_{h}(p-w)\|_{L^{2}(T)}$ as done in 117. If $z\in\mathcal{N}_{h}\backslash\partial\Omega$, we use the Cauchy–Schwarz inequality and the fact that $\|\varphi_{z}\|_{L^{2}(T)}\lesssim h_{T}^{n/2}$ to deduce
@@
As suggested by the proof of Lemma 4.3, one can not guarantee that $\mathcal{C}_{h}p(z)=p(z)$ for all $p\in\mathbb{P}_{1}(\omega_{z})$ if 123 does not hold. Thus, a lack of local symmetry results in the proposed quasi-interpolant lacking optimality. Of course, one may be inspired by the construction of $\mathcal{C}_{h}$ in Section 4.1 to generalize the proposed operator 122. In particular, one can define a convex combination of weights $\{\alpha_{z^{\prime}}\}_{z^{\prime}\in\omega_{z}}$ such that the following reweighted operator,
@@
preserves linear functions on the super-macro element $\cup_{z^{\prime}\in\mathcal{N}_{h}\cap\omega_{z}}\omega_{z^{\prime}}$. This construction requires a more delicate analysis and specific assumptions on $\mathcal{T}_{h}$ near the boundary $\partial\Omega$ to guarantee non-negativity of $\mathcal{C}_{h}(v_{h}-\phi)$ for all $v_{h}\in K_{h}\cup K$. Although we save the technical details for future work, we perform this type of construction on a one-dimensional subdomain for the Signorini problem considered in Section 5; in particular, see 147.
@@ 4.3. Error estimates for the obstacle problem
[…]
@@
[…]
@@
[…]
@@
(129) $\mathcal{E}_{h}w:=\mathcal{C}_{h}(w-\phi)+\phi+\epsilon_{h}.$
where $\mathcal{C}_{h}$ is the appropriate quasi-interpolant analyzed in Lemmas 4.2 and 4.3 and
(130) $\epsilon_{h}=\sum_{z\in\mathcal{N}_{h}\backslash\partial\Omega}\epsilon\varphi_{z},\quad\epsilon=\mathrm{card}(\mathcal{T}_{h})^{-1/2}\min_{T\in\mathcal{T}_{h}}h_{T}^{2-n/2}.$
Observe that $\mathcal{E}_{h}w|_{\partial\Omega}=0$ for any $w\in H^{1}_{0}(\Omega)$, which follows from $\mathcal{C}_{h}w|_{\partial\Omega}=\epsilon_{h}|_{\partial\Omega}=0$ and $(\phi-\mathcal{C}_{h}\phi)|_{\partial\Omega}=0$ since $\phi$ is a constant on $\partial\Omega$. Hence, we obtain that $\mathcal{E}_{h}w\in H^{1}_{0}(\Omega)$ because $\mathcal{C}_{h}(w-\phi)+\epsilon_{h}\in\mathbb{P}_{1}(\mathcal{T}_{h})\cap H^{1}(\Omega)$ and $\phi\in H^{1}(\Omega)$.
We now verify that $\operatorname{ess\,inf}(\mathcal{E}_{h}w-\phi)>0$ in $\Omega$ whenever $w\in K\cup K_{h}$. This is necessary to ensure that $\mathcal{E}_{h}w\in\operatorname{dom}(\nabla\mathcal{R})$ where $\nabla\mathcal{R}=\ln(u-\phi)$ for Algorithm 2, see 97. By construction of the interpolants, cf. 114 and 122, we have that $\mathcal{C}_{h}(w-\phi)(z)\geq 0$ for any $z\in\mathcal{N}_{h}$. In addition, for $z\in\partial\Omega$, $\mathcal{C}_{h}(w-\phi)(z)=-\phi|_{\partial\Omega}=\delta$ since $w|_{\partial\Omega}=0$. Therefore,
$\mathcal{C}_{h}(w-\phi)+\epsilon_{h}\geq\sum_{z\in\partial\Omega}\delta\varphi_{z}+\sum_{z\in\mathcal{N}_{h}\backslash\partial\Omega}\epsilon\varphi_{z}\geq\min(\delta,\epsilon)\sum_{z\in\mathcal{N}_{h}}\varphi_{z}=\min(\delta,\epsilon)>0.$
Hence, we conclude that $\mathcal{E}_{h}w-\phi=\mathcal{C}_{h}(w-\phi)+\epsilon_{h}\geq\min(\delta,\epsilon)>0$.
Observing that $\mathcal{E}_{h}w=\mathcal{C}_{h}w+\varepsilon$ where $\varepsilon=\phi-\mathcal{C}_{h}\phi+\epsilon_{h}$, the inequality 85b is verified in Lemma 4.2 and Lemma 4.3. It remains to verify 85c. Using the triangle inequality followed by 90, we find that
(131) $\|\varepsilon\|_{H^{t}(\Omega)}$ $\leq\|\phi-\mathcal{C}_{h}\phi\|_{H^{t}(\Omega)}+\|\epsilon_{h}\|_{H^{t}(\Omega)}$
$\lesssim h^{1+s-t}|\phi|_{H^{1+s-t}(\Omega)}+\|\epsilon_{h}\|_{H^{t}(\Omega)}.$
Next, we estimate $\|\epsilon_{h}\|_{H^{t}(\Omega)}$. To this end, observe that $\|\phi_{z}\|_{L^{2}(T)}\lesssim h_{T}^{n/2}$ and $\epsilon_{h}|_{T}=\sum_{z\in\mathcal{V}_{T}\backslash\partial\Omega}\epsilon\varphi_{z}$ for any $T\in\mathcal{T}_{h}$. Therefore, with the fractional inverse inequality (see e.g., [7, Proposition 4.2.12]),
(132) $\|\epsilon_{h}\|_{H^{t}(T)}^{2}\lesssim h_{T}^{-2t}\|\epsilon_{h}\|^{2}_{L^{2}(T)}\lesssim\sum_{z\in\mathcal{V}_{T}\backslash\partial\Omega}\epsilon^{2}h_{T}^{-2t}\|\varphi_{z}\|^{2}_{L^{2}(T)}\lesssim\epsilon^{2}h_{T}^{n-2t}\lesssim h_{T}^{4-2t}\mathrm{card}(\mathcal{T}_{h})^{-1}.$
Summing over all elements $T\in\mathcal{T}_{h}$ shows that $\|\epsilon_{h}\|_{H^{t}(\Omega)}\lesssim h^{2-t}\lesssim h^{1+s-t}$ since $s\in[0,1]$. This bound and 131 yield the required estimate 85c. Having verified both Assumptions 2 and 3, we invoke Theorem 3.8 to conclude that 127 holds.
To show 128, we utilize Lemma 3.6 and note that $P_{W_{h}}\tilde{u}_{h}^{\ell}=\tilde{u}_{h}^{\ell}$ since $\tilde{u}_{h}^{\ell}\in\mathbb{P}_{0}(\mathcal{T}_{h})$. Therefore, since $B$ is the identity map and $\Omega_{d}=\Omega$ in this case, 79 reads
@@
where we used the stability of $P_{W_{h}}$ and that $\|u^{*}-P_{W_{h}}u^{*}\|_{L^{2}(T)}\lesssim h_{K}|u^{*}|_{H^{1}(T)}$ for any $T\in\mathcal{T}_{h}$. The final estimate 128 is obtained by combining 127 and the final inequality in 133. ∎

### v2

@@ 4. Application I: The obstacle problem
and the obstacle $\phi\in H^{1}(\Omega)\cap L^{\infty}(\Omega)$ is fixed. For simplicity, we assume that $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\phi|_{\partial\Omega}=-\delta\leq 0$} for some non–negative constant $\delta$. It is well-known that there exists a unique solution to 87 [16, Theorem 6.9-1], here denoted by $u^{\ast}$.
@@
(91a) $\alpha_{k}\,(\nabla u^{k}_{h},\nabla v_{h})+(v_{h},\psi^{k}_{h}-\psi_{h}^{k-1})$ $=\alpha_{k}(f,v_{h})$  $~\text{ for all }v_{h}\in V_{h},$
(91b) $(u^{k}_{h},w_{h})-(\exp(\psi^{k}_{h})+\phi,w_{h})$ $=0$  $~\text{ for all }w_{h}\in W_{h}.$
@@
Main goal: We derive error estimates in Corollary 2 for the PG method applied to the obstacle problem (Algorithm 2) for the two choices of $V_{h}\times W_{h}$ given in 94 and 95, below. To this end, we utilize Theorem 3.6 and verify that Assumptions 2 and 3 hold for each pair of spaces, respectively. We follow with preliminary results for each case.
Case I. $(\mathbb{P}_{1}\text{-bubble},\mathbb{P}_{0}\text{-broken})$ Define the space
@@
(94) $V_{h}$ $=\{v\in L^{\infty}(\Omega)\mid v|_{T}\in\mathbb{P}_{1}(T)\oplus\mathbb{B}(T)~\text{ for all }T\in\mathcal{T}_{h}\}\cap H^{1}_{0}(\Omega),$
@@
Both of the $V_{h}$ subspaces in 94 and 95 imply fixing the degrees of freedom of $u_{h}^{k}$ on the Dirichlet boundary. This is a textbook procedure that is simplified in our setting because of the homogeneous boundary conditions specified in 87. Likewise, the continuous Lagrange elements defining $W_{h}$ in 95 also require fixing the degrees of freedom of the latent variable $\psi_{h}^{k}$ on the Dirichlet boundary. In this case, we construct the lift
@@
where $\mathcal{N}_{h}\cap\partial\Omega$ are the boundary dofs and $\{\varphi_{z}\}_{z\in\mathcal{N}_{h}}$ are the global shape functions corresponding to the (nodal) dofs $\mathcal{N}_{h}$. This construction ensures that the bound-preserving discrete solution $\tilde{u}_{h}^{k}$ 92 satisfies the same homogeneous Dirichlet boundary conditions as the solution $u^{\ast}$ it’s meant to approximate. Indeed, observe that
@@
with $\nabla\mathcal{R}^{*}(\psi)=\exp(\psi)+\phi$ and $\nabla\mathcal{R}(u)=\ln(u-\phi)$ for the particular choice of Legendre function in Algorithm 2. In turn, when using the elements in Case II, we understand 91b as
@@
4.1. Preliminaries for the $(\mathbb{P}_{1}\text{-bubble},\mathbb{P}_{0}\text{-broken})$-element pair
We begin with the Fortin operator for the subspaces in 94, and establish the operator’s approximation properties required in Assumption 3.
@@ 4.1. Preliminaries for the $(\mathbb{P}_{1}\text{-bubble},\mathbb{P}_{0}\text{-broken})$-element pair
The subspaces given in 94 satisfies the inf-sup condition 40 and there exists a stable Fortin operator satisfying 42 and
@@
where $\tilde{\mathcal{I}}_{h}:L^{1}(\Omega)\rightarrow\mathbb{P}_{1}(\mathcal{T}_{h})\cap H^{1}_{0}(\Omega)$ is the quasi-interpolant introduced in [23, Section 6], $I$ is the identity operator, and $\tilde{\Pi}_{h}:L^{2}(\Omega)\rightarrow V_{h}$ is defined element-wise to satisfy $(\tilde{\Pi}_{h}v)|_{T}:=b_{T}v_{T}\in\mathbb{B}(T)$, where $v_{T}\in\mathbb{P}_{0}(T)=\mathbb{R}$ solves
@@
Using 102 in 99 along with the triangle inequality and the stability of $\tilde{\mathcal{I}}_{h}$, we obtain the stability of $\Pi_{h}$ in $L^{2}(\Omega)$. To show stability in the $H^{1}$-semi norm, we apply triangle inequality, a local inverse estimate, 102 and the stability and approximation properties of $\tilde{\mathcal{I}}_{h}$ [23, Lemma 6.3 and Theorem 6.4]:
@@
where $\Delta_{T}$ is a macro-element. Summing over mesh elements and using mesh regularity shows stability in the $H^{1}(\Omega)$-seminorm. We conclude that 97 holds. This stability estimate and the observation that
@@
Note that the discrete iterates $u_{h}^{k}$ have bound-preserving local averages. Indeed, testing 91b by the indicator function of one element (an admissible test function because $W_{h}=\mathbb{P}_{0}(\mathcal{T}_{h})$) and using that $\nabla\mathcal{R}^{*}(\psi)>\phi$ shows that $u_{h}^{k}\in K_{h}$ for all $k$, with
@@
To satisfy Assumption 2, we first require a map defined over $V$ and mapping the set $K_{h}$ to the set $K$. We will later shift this map to construct the enriching map $\mathcal{E}_{h}$, see 121. The natural choice is the Clément interpolant [17, 5]. Here, we utilize a specialized variant of this interpolant that employs weighted local averages over vertex patches to maintain second-order accuracy for smooth functions. This specific interpolant was first introduced in [29], but we slightly modify its definition to incorporate non-homogeneous boundary data.
@@
(107) $\|w-\mathcal{C}_{h}w\|_{L^{2}(T)}+h_{T}\|\nabla(w-\mathcal{C}_{h}w)\|_{L^{2}(T)}$ $\lesssim h_{T}^{1+s}|w|_{H^{1+s}(\Delta_{T})}$  $~\text{ for all }w\in H^{1+s}(\Delta_{T}),$
@@
In the above, we also used that $\mathrm{card}(\omega_{z})$ is uniformly bounded w.r.t. $h$ for all $z\in\mathcal{N}_{h}$, which follows from the shape-regularity of the mesh; see, e.g., [24, Proposition 11.6]. If $z\in\partial\Omega$, then $z\in\partial T\cap\omega_{z}$ for some $T\subset\omega_{z}$. Using the definition of $\mathcal{C}_{h}$ on boundary nodes, we apply local inverse and trace inequalities to bound
@@
Combining 109 with 112 and using the Bramble–Hilbert Lemma (see, e.g., [23, Lemma 5.6]) yields the required bound on the first term of 107. To obtain the required bound on the second term, we apply triangle and inverse estimates:
@@ 4.2. Preliminaries for the $(\mathbb{P}_{1},\mathbb{P}_{1})$-element pair
For this case, the Fortin operator is simply the $L^{2}(\Omega)$-projection; cf. [36, Remark 5.2]. Thus, the main task is to construct the enriching map of Assumption 2, which will again depend on a modified Clément quasi-interpolant. Using the definition of $K_{h}$ 74 and testing with $\varphi_{z}$, we observe that
@@
If $z\in\partial\Omega$, we proceed as before and select $v_{z}=(\mathcal{SZ}_{h}v)(z)$. For simplicity, we make the following symmetry assumption on the mesh, which is sufficient for optimality of the quasi-interpolant given in 114. Note that this assumption is the same condition necessary for optimality of the classical Clément quasi-interpolant [29].
@@
We now prove that this condition implies optimality of 114.
@@
(116) $\|w-\mathcal{C}_{h}w\|_{L^{2}(T)}+h_{T}\|\nabla(w-\mathcal{C}_{h}w)\|_{L^{2}(T)}$ $\lesssim h_{T}^{1+s}|w|_{H^{1+s}(\Delta_{T})}$  $~\text{ for all }w\in H^{1+s}(\Delta_{T}),$
@@
The proof follows similar arguments to the proof of Lemma 4.2. We only highlight the key points. We first show that under 115, $\mathcal{C}_{h}p(z)=p(z)$ for all $p\in\mathbb{P}_{1}(\omega_{z})$. To see this first note that for linear $p$, we have
@@
where we used that $\int_{T}\varphi_{z^{\prime}}\varphi_{z}\operatorname{d}\!x=(n+1)^{-1}(n+2)^{-1}|T|(1+\delta_{z^{\prime},z})$ [15, Exercise 4.1.1]. Hence,
@@
In the last step, we used that $\int_{\omega_{z}}\varphi_{z}=|\omega_{z}|/(n+1)$ and 115. Along with the observation that $\overline{b}_{z}=b$ for any constant $b$, we conclude that $\mathcal{C}_{h}p=p$ for any $p\in\mathbb{P}_{1}(\omega_{z})$. In turn, it suffices to bound $\|\mathcal{C}_{h}(p-w)\|_{L^{2}(T)}$ as done in 109. If $z\in\mathcal{N}_{h}\backslash\partial\Omega$, we use the Cauchy–Schwarz inequality and the fact that $\|\varphi_{z}\|_{L^{2}(T)}\lesssim h_{T}^{n/2}$ to deduce
@@
As suggested by the proof of Lemma 4.3, one can not guarantee that $\mathcal{C}_{h}p(z)=p(z)$ for all $p\in\mathbb{P}_{1}(\omega_{z})$ if 115 does not hold. Thus, a lack of local symmetry results in the proposed quasi-interpolant lacking optimality. Of course, one may be inspired by the construction of $\mathcal{C}_{h}$ in Section 4.1 to generalize the proposed operator 114. In particular, one can define a convex combination of weights $\{\alpha_{z^{\prime}}\}_{z^{\prime}\in\omega_{z}}$ such that the following reweighted operator,
@@
preserves linear functions on the super-macro element $\cup_{z^{\prime}\in\mathcal{N}_{h}\cap\omega_{z}}\omega_{z^{\prime}}$. This construction requires a more delicate analysis and specific assumptions on $\mathcal{T}_{h}$ near the boundary $\partial\Omega$ to guarantee non-negativity of $\mathcal{C}_{h}(v_{h}-\phi)$ for all $v_{h}\in K_{h}\cup K$. Although we save the technical details for future work, we perform this type of construction on a one-dimensional subdomain for the Signorini problem considered in Section 5; in particular, see 137.
@@ 4.3. Error estimates for the obstacle problem
For both the $(\mathbb{P}_{1}\text{-bubble},\mathbb{P}_{0}\text{-broken})$ and $(\mathbb{P}_{1},\mathbb{P}_{1})$ element pairs, the following error estimate holds:
@@
In addition, for the $(\mathbb{P}_{1}\text{-bubble},\mathbb{P}_{0}\text{-broken})$ elements, we have the following estimate on the bound-preserving approximation $\tilde{u}_{h}^{\ell}$:
@@
We verify the assumptions of Theorem 3.6. For the $(\mathbb{P}_{1}\text{-bubble},\mathbb{P}_{0}\text{-broken})$ elements, Lemma 4.1 shows that 76a-76b and 78a in Assumption 3 hold. For the $(\mathbb{P}_{1},\mathbb{P}_{1})$ elements, the Fortin operator is simply the $L^{2}(\Omega)$ projection onto $V_{h}$. It is standard to show that 76a-76b and 78a hold if the mesh is quasi-uniform, as assumed in this case; see, e.g., [24, Proposition 22.21] and [36, Remark 5.2].
@@
(121) $\mathcal{E}_{h}w:=\mathcal{C}_{h}(w-\phi)+\phi$
where $\mathcal{C}_{h}$ is the appropriate quasi-interpolant analyzed in Lemmas 4.2 and 4.3. Observe that $\mathcal{E}_{h}w|_{\partial\Omega}=0$ for any $w\in H^{1}_{0}(\Omega)$, which follows from $\mathcal{C}_{h}w|_{\partial\Omega}=0$ and $(\phi-\mathcal{C}_{h}\phi)|_{\partial\Omega}=0$ since $\phi$ is a constant on $\partial\Omega$. Hence, we obtain that $\mathcal{E}_{h}w\in H^{1}_{0}(\Omega)$ because $\mathcal{C}_{h}(w-\phi)\in\mathbb{P}_{1}(\mathcal{T}_{h})\cap H^{1}(\Omega)$ and $\phi\in H^{1}(\Omega)$. By construction of the interpolants, cf. 106 and 114, we have that $\mathcal{C}_{h}(w-\phi)(z)\geq 0$ for any $z\in\mathcal{N}_{h}$. In addition, for $z\in\partial\Omega$, $\mathcal{C}_{h}(w-\phi)(z)=-\phi|_{\partial\Omega}\geq 0$. Hence, we conclude that $\mathcal{E}_{h}w\geq\phi$ and $\mathcal{E}_{h}w\in K$.
Observing that $\mathcal{E}_{h}w=\mathcal{C}_{h}w+\varepsilon$ where $\varepsilon=\phi-\mathcal{C}_{h}\phi$, the inequality 78b is verified in Lemma 4.2 and Lemma 4.3. It remains to verify 78c. This follows directly from 83
(122) $\|\varepsilon\|_{H^{t}(\Omega)}$ $=\|\phi-\mathcal{C}_{h}\phi\|_{H^{t}(\Omega)}\lesssim h^{1+s-t}|\phi|_{H^{1+s-t}(\Omega)}.$
Having verified both Assumptions 2 and 3, we invoke Theorem 3.6 to conclude that 119 holds.
To show 120, we utilize Lemma 3.5 and note that $P_{W_{h}}\tilde{u}_{h}^{\ell}=\tilde{u}_{h}^{\ell}$ since $\tilde{u}_{h}^{\ell}\in\mathbb{P}_{0}(\mathcal{T}_{h})$. Therefore, since $B$ is the identity map and $\Omega_{d}=\Omega$ in this case, 71 reads
@@
where we used the stability of $P_{W_{h}}$ and that $\|u^{*}-P_{W_{h}}u^{*}\|_{L^{2}(T)}\lesssim h_{K}|u^{*}|_{H^{1}(T)}$ for any $T\in\mathcal{T}_{h}$. The final estimate 120 is obtained by combining 119 and the final inequality in 123. ∎